# On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems

### v1

@@ 5 Theoretical Analysis
Notations. The dataset of the target domain is $D_{T}$ and the input data to the model is $\mathbf{x}$ and the output data is $\mathbf{y}$, which the model should predict given the input data. There are $N$ number of DAFT models, and a DAFT model can be built using any base model $B_{j}\in\mathcal{B}$, with $j=\{1,2,...,J\}$ and fine-tuning it on dataset $D_{k}\in\mathcal{DA}$ with $k=\{1,2,...,K\}$, where $\mathcal{DA}$ is the set of domain adjacent datasets. The $i^{th}$ DAFT model is given by $M_{i}$ and the corresponding base model and fine-tuning dataset are given by $B_{\kappa(i)}$ and $D_{\nu(i)}$, respectively, i.e., $B_{\kappa(i)}$ was fine-tuned on $D_{\nu(i)}$ dataset to get $M_{i}$. Here, $\kappa(i)$ and $\nu(i)$ are index mapping functions. Also, if we assume that all these DAFT models were created by fine-tuning fully on the given datasets (no fractional fine-tuning), then the total number of DAFT can not be greater than $J\times K$, or $N\leq JK$. Lastly, let $\ell(M(\mathbf{x}),\mathbf{y})$ be the loss in output when model $M$ is used on target dataset input $\mathbf{x}$ and the output is compared with $\mathbf{y}$. Since $\mathbf{x}$ and $\mathbf{y}$ are common for all loss calculation, we can just use $\ell(M)$ to represent $\ell(M(\mathbf{x}),\mathbf{y})$.
5.1 DAFT-E ${}^{\text{Z}}$ versus DAFT ${}^{\text{Z}}$
It is hard to know if a DAFT model is going to perform well or not a priori, and instead of choosing one of the DAFT models in random (DAFT ${}^{\text{Z}}$), one solution could be to use the ensemble of the output results (DAFT-E ${}^{\text{Z}}$) from all the DAFT models.
@@ 5.1 DAFT-E ${}^{\text{Z}}$ versus DAFT ${}^{\text{Z}}$
The performance of the average ensemble of DAFT models (DAFT-E ${}^{\text{Z}}$) is no worse than the expected performance obtained from choosing the DAFT models uniformly at random.
Proposition 1 states that in terms of expected performance, using DAFT-E ${}^{\text{Z}}$ is no worse (strictly better, if $\ell$ is assumed to be strictly convex) than picking from the DAFT models uniformly at random.
@@ 5.2 DAFT-E versus Optimum
[…]
$\;+\min_{i\in\mathcal{N}}\left[\mu(\tilde{M}_{i},\tilde{M_{*}})+\rho(D_{\nu(i)},D_{T})\right],$ (1)
where $\mu(\tilde{M}_{i},\tilde{M_{*}})$ is the performance difference between the models $\tilde{M}_{i}$ and $\tilde{M_{*}}$ both fine-tuned on $D_{T}$, and $\rho(D_{\nu(i)},D_{T})$ is an appropriately defined distance measure between $D_{\nu(i)}$ and $D_{T}$.
@@
Proposition 2 and Corollary 1 (which follows directly from Proposition 2) imply that the performance of DAFT-E depends on the base model on which the DAFT models were trained and the datasets they were fine-tuned on. If the base models (FMs) in the ensemble are nearly as good as the best possible base model (FM) (when the performance of the FFT of the base models are compared), then we can ignore the $\mu()$ term in Equation 1, and the performance of DAFT-E depends only on the dataset that is closest to the target domain. If the DAFT models are generated from a large number of diverse domain adjacent datasets, we expect the distance of the target dataset to the closest DAFT dataset to be small, resulting in DAFT-E performing very close to the optimum.
@@ 6 Conclusion
In this paper we introduce the notion of domain adjacent fine-tuned models (DAFT), and investigate how DAFT models can be utilized for inference under limitations on computation time or data required for training a model on a target task. The DAFT models can be used on the target domain (zero-shot) without any fine-tuning. If chosen properly, the performance of DAFT models on the target domain can be close to the optimal performance. However, this performance depends strongly on the choice of the DAFT model, and ensemble methods can be used to address this issue. Two ensembling methods, DAFT-E ${}^{\text{Z}}$ and DAFT-E are explored, and their performances are empirically evaluated and compared against individual DAFT models and other benchmarks involving larger LLMs and base models trained with the full target domain dataset. Our theoretical results support the conclusions from the empirical findings, and provide insights under what conditions DAFT-E can provide near-optimal performance.
7 Limitations
Here in this section we discuss a few limitations and areas of improvement for using DAFT models or ensembles of them.
  • Non domain-adjacent model: DAFT-E or DAFT-E ${}^{\text{Z}}$ will not perform well when none of the available DAFT models are adjacent to the target domain.
  • Sparse weight in Ensemble Layer: Currently DAFT-E ${}^{\text{Z}}$ and DAFT-E needs inference results from all the DAFT models; a sparse weight ensemble may help us select a subset of the DAFT models to have efficient implementation of DAFT-E.
  • Availability of DAFT models: In the ensemble methods discussed, the DAFT models are not trained, rather are used directly from off the shelf in the simplest way. Our argument is that new DAFT models are being generated, and with time DAFT models will be much easier to find. However that is not yet true for some of the tasks being currently investigated in the literature.
  • Run Time: Upon running the DAFT-E and FT in the CPU, we found that the runtime of FT is 6 times more than the inference of the same sample size (Roberta model). Hence, a DAFT-E that ensembles more than 6 DAFT models will be slower than FT, unless the execution of the DAFT models are done in parallel.
@@ A.1 Datasets and Models
The attributes of the six datasets used for sentiment analysis and the three dataset used for text similarity analysis are given in Table 4 and 5. For ‘SST2’ we had to use validation dataset as test data, because the test data does not have any labels. The direct link to download these datasets are given as follows: https:// huggingface.co /datasets/amazon_polarity, ‘https:// www.cs.cornell.edu /people/pabo/movie-review-data/’, ‘https://huggingface.co/datasets/ imdb’, ‘https:// huggingface.co/ datasets/sst2’, ‘https:// huggingface.co/ datasets/ mteb/tweet_sentiment_extraction’, ‘https:// huggingface.co/ datasets/ yelp_polarity’, ’https://huggingface.co /datasets/nyu-mll/glue’ (when copying the links please remove any spaces from the texts). Figure 6 and 7 shows the performance heat-map for sentiment analysis task when the base models are BERT-based-uncased and xlnet-base-cased respectively. The heat maps in general follows a similar pattern as observed in Figure 2, i.e., Tweet test data cannot be predicted well by any other DAFT models, and Cornell test data is also hard to predict. The observation of DAFT trained on tweet data performing poor is true for bert-based-uncased as well, but not always true for xlnet-base-cased. Another interesting observation is that for xlnet, the DAFT model trained on IMDB data seems to perform quite poorly in a few instances (i.e., on test data of Cornell and Tweet).
@@ A.2 Experimental Settings
A.3 Discussion on Performance Anomalies
A.3.1 SST2 on Heatmap
The SST2 dataset obtained from Huggingface did not have any label on its test data, and we had to use the validation data of SST2 as the test data instead. Also, the validation data of SST2 only had $872$ samples. Now, in the heat map shown in Figure 2, the performance of DAFT-SST2 achieved $92\%$ accuracy on the SST2 validation data, whereas, DAFT-Cornell achieved $93.1\%$ accuracy on that same test dataset. This is quite counter intuitive, because we expect DAFT-SST2 to be the best performing one on its own domain (SST2 validation data). We did a thorough simulation analysis to check what might be the reason. Since the DAFT-SST2 was generated using randomly chosen $7,000$ samples from the training dataset of SST2, and we initially extracted the rest of the training data and divided it to different batches of $5,000$ samples. Now, with this batches we checked the performance of DAFT- SST2 and DAFT- Cornell. As expected, we got better performance with DAFT- SST2 compared to DAFT- Cornell with these new batches. With this experiment, we concluded that SST2 validation dataset is, for some unknown reason, more similar to the Cornell training dataset than the SST2 training data, and thus giving a counter intuitive result.
A.3.2 STSB with DAFT-E
The performance of DAFT-E with $LR$ suffers a lot compared to both DA(FT)2 that have been considered here, i.e., DAFT-mrpc and DAFT-qqp. Both of these DAFT models have roberta as their base model, and interestingly these roberta based DAFT models seemed to perform well even at zero shot on STSB dataset. Whereas, the DAFT models based on BERT and xlnet showed considerable less performance. We found this to be the main reason for which DAFT-E performed poorly in comparison to the DA(FT)2 models which already had better performance even without any fine-tuning. To check this, we even checked the DAFT-E performance with DAFT-mrpc and DAFT-qqp with roberta base only, and found that DAFT-E outperformed the DA(FT)2 up to $n=64$.
@@ A.4 Fine-tuning the header or whole model
Figure 10 to 13 show the performance comparison of fine-tuning only the header layer of the base model and fine-tuning the whole model of the base models. It is straightforward to see the superior performance of fine-tuning the whole model even though the base models (i.e., BERT, Roberta) have already gone through extensive training with language data.
@@
A.5 Ensemble Layer of DAFT-E
A.5.1 Weight update of $LR$
[…]
@@ A.5.1 Weight update of $LR$
A.5.2 Parameters of $LR$ and $RF$
@@
A.6 Performance of DAFT models and DAFT-E ${}^{\text{Z}}$ for Textual similarity task
Figure 14 and 15 show the Relative Improvement of DAFTs compared to Single Best DAFT and DAFT-E ${}^{\text{Z}}$ respectively for the textual similarity task with three datasets. Similar to the results with sentiment analysis task, we observe the performance deviation among the DAFT models and usual better performance of DAFT-E ${}^{\text{Z}}$.
Figure 14: Relative Improvement of DAFTs compared to the Single Best DAFT for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the single-best DAFT FM (out of the 6). Values less than 0 indicate performance degradation.
Figure 15: Relative Improvement of DAFTs compared to DAFT-E ${}^{\text{Z}}$ for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the DAFT-E ${}^{\text{Z}}$. Values less than 0 indicate performance degradation.
A.7 Proof of Propositions

### v2

@@ 5 Theoretical Analysis
Notations. The dataset of the target domain is $D_{T}$, the input data to the model is $\mathbf{x}$ and the output data is $\mathbf{y}$, which the model should predict given the input data. There are $N$ number of DAFT models, and a DAFT model can be built using any base model $B_{j}\in\mathcal{B}$, with $j=\{1,2,...,J\}$ and fine-tuning it on dataset $D_{k}\in\mathcal{DA}$ with $k=\{1,2,...,K\}$, where $\mathcal{DA}$ is the set of domain adjacent datasets. The $i^{th}$ DAFT model is given by $M_{i}$ and the corresponding base model and fine-tuning dataset are given by $B_{\kappa(i)}$ and $D_{\nu(i)}$, respectively, i.e., $B_{\kappa(i)}$ was fine-tuned on $D_{\nu(i)}$ dataset to get $M_{i}$. Here, $\kappa(i)$ and $\nu(i)$ are index mapping functions. Also, if we assume that all these DAFT models were created by fine-tuning fully on the given datasets (no fractional fine-tuning), then the total number of DAFT can not be greater than $J\times K$, or $N\leq JK$. Lastly, let $\ell(M(\mathbf{x}),\mathbf{y})$ be the loss in output when model $M$ is used on target dataset input $\mathbf{x}$ and the output is compared with $\mathbf{y}$. Since $\mathbf{x}$ and $\mathbf{y}$ are common for all loss calculation, we can just use $\ell(M)$ to represent $\ell(M(\mathbf{x}),\mathbf{y})$.
5.1 DAFT-E ${}^{\text{\!\scriptsize Z}}$ versus DAFT ${}^{\text{\!\scriptsize Z}}$
Determining if a DAFT model is going to perform well or not is very hard to know a priori if adequate time or data to train or test the model is not available. Hence, instead of choosing one of the DAFT models in random (DAFT ${}^{\text{\!\scriptsize Z}}$), one solution could be to use the ensemble of the output results (DAFT-E ${}^{\text{\!\scriptsize Z}}$) from all the DAFT models.
@@ 5.1 DAFT-E ${}^{\text{\!\scriptsize Z}}$ versus DAFT ${}^{\text{\!\scriptsize Z}}$
The performance of the average ensemble of DAFT models (DAFT-E ${}^{\text{\!\scriptsize Z}}$) is no worse than the expected performance obtained from choosing the DAFT models uniformly at random.
Proposition 1 states that in terms of expected performance, using DAFT-E ${}^{\text{\!\scriptsize Z}}$ is no worse (strictly better, if $\ell$ is assumed to be strictly convex) than picking from the DAFT models uniformly at random. From the performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$ shown in Fig. 5 and 6, we see that Proposition 1 holds, i.e., DAFT-E ${}^{\text{\!\scriptsize Z}}$ has better performance than the average performance of the DAFT models. Moreover, for these two specific types of tasks, the ensemble of the DAFT models beats the performance of individual DAFT models in most cases (98 out of 108 in total for both tasks).
@@ 5.2 DAFT-E versus Optimum
where $\mu(\tilde{M}_{i},\tilde{M_{*}})$ is the performance difference between the models $\tilde{M}_{i}$ and $\tilde{M_{*}}$ both fine-tuned on $D_{T}$, and $\rho(D_{\nu(i)},D_{T})$ [^12] Some practical dataset distance measurement metric are the Wasserstein distance metric (Panaretos & Zemel, 2019), Jensen-Shannon Divergence (JSD), etc. is an appropriately defined distance measure between $D_{\nu(i)}$ and $D_{T}$.
Usually base models (derived from FMs which are all quite large) perform similarly when fine-tuned on the full target dataset. Under that assumption, we can ignore the $\mu()$ term in the bound, resulting in the following corollary.
@@
Proposition 2 and Corollary 1 imply that the performance of DAFT-E depends on the base model on which the DAFT models were trained and the datasets they were fine-tuned on. If the base models (FMs) in the ensemble are nearly as good as the best possible base model (FM) (when the performance of the FFT of the base models are compared), then we can ignore the $\mu()$ term in Equation 1, and the performance of DAFT-E depends only on the dataset that is closest to the target domain. If the DAFT models are generated from a large number of adjacent datasets from diverse domains, we expect the distance from the target dataset to the closest DAFT dataset to be small, resulting in DAFT-E performing very close to the optimum.
In our experimental setting, the assumption made for the Corollary 1 holds, that is, all three base models are comparable in terms of performance when fine-tuned on the same dataset. Hence, the performance of DAFT-E depends on the adjacency of the datasets used in the DAFT models with the target dataset. For the case of sentiment analysis, we have six datasets, whereas for the text similarity task, we only have three. Thus, it is more probable that the DAFT models of the text similarity task are more diverse and DAFT-E will be able to achieve close to optimum performance. From the results (Fig. 8, and 9) it is evident that DAFT-E was able to perform much better for the sentiment analysis task, compared to the textual similarity task.
@@ 6 Conclusion
In this paper, we explore 4 ways to leverage the abundance of publicly available fine-tuned LLMs for data-scarce tasks. We investigate how the DAFT models can be utilized for inference under limitations on computation time or data required for training a model on a target task. The DAFT models can be used on the target domain (zero-shot) without any fine-tuning. If chosen properly, the performance of DAFT models on the target domain can be close to the optimal performance. However, this performance depends strongly on the choice of the DAFT model, and ensemble methods can be used to address this issue. Two ensemble methods, DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E are explored; their performances are empirically evaluated and compared against individual DAFT models and other benchmarks involving larger LLMs and base models trained with the full target domain dataset. Our theoretical results support the conclusions from the empirical findings, and provide insights under what conditions DAFT-E can provide near-optimal performance.
Broader Impact Statement
The use of DAFT models can be a great low-resource and easy-to-use solution in data-scarce domains. Due to the current surge of open source DAFT models for different downstream tasks, the community can use these models directly without any (computationally expensive) fine-tuning of LLMs. The main innovation of the proposed ensemble methods, i.e., DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E, is the utilization of readily available DAFT models to perform lightweight solutions for problems where little or no data from the target domain is available.
On the other hand, there are certain limitations of our method that the users may face and should keep in mind. Firstly, the performance of these methods is highly dependent on the availability of DAFT models. DAFT-E or DAFT-E ${}^{\text{\!\scriptsize Z}}$ will not perform well when none of the available DAFT models are adjacent to the target domain. Moreover, while many DAFT models for sentiment analysis and NLI tasks are currently available, DAFTs for some other tasks might not be as widely available soon. Secondly, DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E require inference with multiple DAFT models. While the multi-model inference can be parallelized, thereby reducing inference time in scenarios where multiple smaller servers are available, inference time would increase if memory is limited or the per-model inference needs to be sequential. Further, with DAFT-E, we can filter out some of the candidate DAFT models, thereby reducing $N$, but that depends on having some data from the target domain. Thirdly, with the proposed methods, the users are supposed to use openly available fine-tuned models; however, the quality or correctness of the fine-tuning datasets and training methods may be questionable. There is also the question of the privacy of the data and the use of the models, which the user might not be aware of. Lastly, the risk of bias and misuse of using the DAFT models is possible. However, in this work, we are not proposing new models, just some methodologies to use already available models through public repositories. Since we are not changing the DAFT models, if the fine-tuning of the DAFT models were done according to safety and bias alignment, we expect to have minimal risk of bias and misuse from using ensembles of these models.
Acknowledgments
The work was supported by RPI-IBM Future of Computing Research Collaboration (https://fcrc.rpi.edu).
@@ A.1 Datasets and Models
The attributes of the six datasets used for sentiment analysis and the three dataset used for text similarity analysis are given in Table 8 and 8. For ‘SST2’ we had to use validation dataset as test data, because the test data does not have any labels. The direct link to download these datasets are given as follows: https:// huggingface.co /datasets/amazon_polarity, ‘https:// www.cs.cornell.edu /people/pabo/movie-review-data/’, ‘https://huggingface.co/datasets/ imdb’, ‘https:// huggingface.co/ datasets/sst2’, ‘https:// huggingface.co/ datasets/ mteb/tweet_sentiment_extraction’, ‘https:// huggingface.co/ datasets/ yelp_polarity’, ’https://huggingface.co /datasets/nyu-mll/glue’ (when copying the links please remove any spaces from the texts). Figs. 12 and 12 shows the performance heat-map for sentiment analysis task when the base models are BERT-based-uncased and xlnet-base-cased respectively. The heat maps in general follows a similar pattern as observed in Fig. 2, i.e., Tweet test data cannot be predicted well by any other DAFT models, and Cornell test data is also hard to predict. The observation of DAFT trained on tweet data performing poor is true for bert-based-uncased as well, but not always true for xlnet-base-cased. Another interesting observation is that for xlnet, the DAFT model trained on IMDB data seems to perform quite poorly in a few instances (i.e., on test data of Cornell and Tweet).
@@ A.2 Experimental Settings
A.3 Discussion on Performance Anomalies - SST2
The SST2 dataset obtained from Huggingface did not have any label on its test data, and we had to use the validation data of SST2 as the test data instead. Also, the validation data of SST2 only had $872$ samples. Now, in the heat map shown in Fig. 2, the performance of DAFT-SST2 achieved $92\%$ accuracy on the SST2 validation data, whereas, DAFT-Cornell achieved $93.1\%$ accuracy on that same test dataset. This is quite counter intuitive, because we expect DAFT-SST2 to be the best performing one on its own domain (SST2 validation data). We did a thorough simulation analysis to check what might be the reason. Since the DAFT-SST2 was generated using randomly chosen $7,000$ samples from the training dataset of SST2, and we initially extracted the rest of the training data and divided it to different batches of $5,000$ samples. Now, with this batches we checked the performance of DAFT- SST2 and DAFT- Cornell. As expected, we got better performance with DAFT- SST2 compared to DAFT- Cornell with these new batches. With this experiment, we concluded that SST2 validation dataset is, for some unknown reason, more similar to the Cornell training dataset than the SST2 training data, and thus giving a counter intuitive result.
@@ A.4 Fine-tuning the header or whole model
Figs. 18 to 18 show the performance comparison of fine-tuning only the header layer of the base model and fine-tuning the whole model of the base models. It is straightforward to see the superior performance of fine-tuning the whole model even though the base models (i.e., BERT, Roberta) have already gone through extensive training with language data.
@@
A.5 Advantage of Ensemble
Our goal in this work is to find a suitable method that ensures better performance when data from the target domain is scarce, and therefore it is not possible to determine conclusively which of the available fine-tuned foundation models are adjacent to the target domain (if DAFT or not) [^13] There are a few established methods (e.g. LEEP score) to find the adjacency between two datasets, however these methods for determining domain-adjacency generally require a substantial amount of data from the target domain.. The advantage of using Ensemble method are: 1) there is no constraint on the DAFT model size and architecture, 2) no need of knowing the weights of the DAFT models, just an API access to the DAFT models is adequate. In our study, we found that even if we do not know how domain-adjacent the available models are, their aggregate ensemble can still perform well (with zero or very limited training) as long as there are some models in the ensemble that are domain-adjacent (although we may not know which ones). Furthermore, as we have observed in the paper, the weighted ensemble of the DAFT models (DAFT-E) can perform similarly to or better than the best single (available) DAFT model. In summary, without adequate data it may be hard to determine which models are DAFT and which are not, and a very low-data low-resource ensembling method like DAFT-E ${}^{\text{\!\scriptsize Z}}$ or DAFT-E can give us good results.
A.6 Ensemble Layer of DAFT-E
A.6.1 Weight update of $LR$
Our ensemble technique uses the final layer output of the DAFT models as the input of the ensemble layer. When performing DAFT-E with $LR$, If $w_{i}$ is the weight of model $i$ in the Ensemble weight layer, then the $LR$ performs the following:
@@ A.6.1 Weight update of $LR$
Figure 19: Relative Improvement of DAFTs compared to the Single Best DAFT for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{\!\scriptsize Z}}$ over the single-best DAFT FM (out of the 6). Values less than 0 indicate performance degradation.
A.6.2 Regression methods: $LR$ and $RF$
We observed the following in terms of performance of the two regression methods: (i) $LR$ has better performance and less variation in performance for smaller sample data, (ii) for higher sample data, i.e., $n$>32, $RF$ usually performs very similar to $LR$, (iii) for sentiment analysis tasks, $LR$ and $RF$ both perform well with similar performance at higher sample data, and (iv) for textual similarity tasks, $LR$ shows minimal performance improvement with more samples, whereas $RF$ shows promising results.
@@ A.6.2 Regression methods: $LR$ and $RF$
A.7 Performance of DAFT models for Textual similarity task
Fig. 19 shows the Relative Improvement of DAFTs compared to Single Best DAFT and DAFT-E ${}^{\text{\!\scriptsize Z}}$ respectively for the textual similarity task with three datasets. Similar to the results with sentiment analysis task, we observe the performance deviation among the DAFT models and usual better performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$.
A.8 Performance of DAFT models from Internet
To evaluate the performance of the readily available DAFT models from public repositories, we have downloaded $100$ sentiment analysis models from Huggingface using ‘popularity’ as the sorting criterion. The performance of these $100$ models compared to the single best DAFT and DAFT-E ${}^{\text{\!\scriptsize Z}}$ is shown in Figs. 21 and 21 respectively. Similar to what we observed previously, DAFT-E ${}^{\text{\!\scriptsize Z}}$ still outperforms most of the DAFT performances ($551$ among $600$ cases). Moreover, it is notable that a lot of the DAFT models are very poor performing compared to our controlled DAFT models introduced in Section 3.
Figure 20: Relative Improvement of DAFTs (from Huggingface) compared to the Single Best DAFT for sentiment analysis task. Values less than 0 indicate performance degradation.
Figure 21: Relative Improvement of DAFTs (from Huggingface) compared to DAFT-E ${}^{\text{\!\scriptsize Z}}$ for sentiment analysis task. Values less than 0 indicate performance degradation.
The performance of DAFT-E for the $100$ models from Huggingface is depicted in Fig. 22. The results are consistent with the results that we observed in Fig. 8. In all cases, there is an upward trend in performance w.r.t. training data with the use of DAFT-E. Also, for $3$ of the $6$ datasets, DAFT-E catches up (or surpasses) the single best DAFT with 128 training pairs. Since we do not have the FFT for this scenario, the FFT comparison is not shown in Fig. 22.
Figure 22: Performance comparison of DAFT-E ${}^{\text{\!\scriptsize Z}}$, DAFT-E, and single-best. Error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble aggregated over 10 trials.
A.9 Proof of Propositions